OpenStudy (anonymous):

a phone company advertises a new plan in which the customer pays a fixed amount of $25 per month for unlimited call in the country,and $0.10 per minute for international calls.Find a rule for the monthly payment a customer pays according to the new plan.write ordered pairs for the monthly payment when the customer uses 90,120,145,and 150 international minutes in a month.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let x = number of minutes spent making international calls

jimthompson5910 (jim_thompson5910):

if it costs $0.10 per minute to make international calls, then how much does it cost if you use up x minutes making international calls ?

OpenStudy (anonymous):

Y=0.10x+25 y=0.10(150)+25 y=40 in other words the user has to pay 40 dallors for using 150 minutes of international calls.

OpenStudy (anonymous):

you just plug the number of minutes in to x

OpenStudy (anonymous):

oo ok thank you @yinkim52001

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

what whould be the ordered pair

OpenStudy (anonymous):

(150,40)

OpenStudy (anonymous):

oo ok

OpenStudy (anonymous):

for one of them

OpenStudy (anonymous):

i got 150,40 because the equation solved for Y and x was the plug in number

OpenStudy (anonymous):

ok i understand thank u for the equation

OpenStudy (anonymous):

no problem